A 90.5-kg ice hockey player hits a 0.125-kg puck, giving the puck a speed of 35 m/s. If both are initially at rest and if the ice is frictionless, how far, in centimeters, does the player recoil in the time it takes the puck to reach the goal, 15 m away?
Answer #1

As initial momentum of system is zero, finally each body will have equal momentum but in opposite direction so that net momentum of system remains zero. (Conservation of linear momentum)

0.0483 m/s

Time taken by puck to travel 15 m

t =distance/speed= 15/35 =(3/7 ) s

Player recoil by

Speed ×time=

0.0483× (3/7)

= 0.0207 m

= 2.07 cm